### v1

@@ Appendix A Proofs of Main Theorem
$\Delta_{3,k}(\widetilde{g}_{i\ell},\ell\neq k)=\frac{1}{T}\sum_{t=1}^{T}f_{it}{\cal E}_{t}\times_{\ell=1,\ell\neq k}^{K}\widetilde{g}_{i\ell}^{\top}\in\mathbb{R}^{d_{k}},$
$\Delta_{4,k}(\widetilde{g}_{i\ell},\ell\neq k)=\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes(f_{jt},\ j\neq i)^{\top}\times_{\ell=1,\ell\neq k}^{K}\widetilde{g}_{i\ell}^{\top}\in\mathbb{R}^{d_{k}\times(r-1)}.$
@@
$\|\Psi\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}\|_{\rm 2}$
$\leq$ $\|\Delta_{1}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{(m)\top}\|_{\rm 2}+\|\Delta_{2,k}(\widehat{g}_{i\ell}^{(m)},\widehat{g}_{i\ell}^{(m)},\ell\neq k)\|_{\rm 2}+2\|\Delta_{3,k}(\widehat{g}_{i\ell}^{(m)},\ell\neq k)\|_{\rm 2}$
@@
$\Delta_{1,1}=a_{ik}\left(\prod_{\ell\neq k}^{K}a_{i\ell}^{\top}\widehat{g}_{i\ell}^{(m)}\right)\frac{1}{T}\sum_{t=1}^{T}f_{i,t}\cdot(f_{j_{2},t},\ j_{2}\neq i)\ {\rm diag}\left(\prod_{\ell\neq k}^{K}a_{j_{2}\ell}^{\top}\widehat{g}_{i\ell}^{(m)},j_{2}\neq i\right)(a_{j_{2}k},j_{2}\neq i)^{\top}.$
@@
Let $\Upsilon_{0,i,k}=T^{-1}\sum_{t=1}^{T}f_{it}^{2}a_{ik}a_{ik}^{\top}$ and $\Upsilon_{0,-i,k}=T^{-1}\sum_{t=1}^{T}\sum_{j_{1},j_{2}\neq i}^{r}f_{j_{1}t}f_{j_{2}t}a_{j_{1}k}a_{j_{2}k}^{\top}$. Then, in the event $\Omega_{1}$,
@@
$\|\Delta_{2,k}\|_{\rm S}\leq\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\right\|_{2}\lesssim\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)}{T}+1,$
$\|\Delta_{3,k}\|_{\rm S}\leq\left\|\frac{1}{T}\sum_{t=1}^{T}f_{it}e_{t}^{\top}\right\|_{2}\lesssim\sqrt{\frac{d\log(d)}{T}}\sqrt{\Theta_{ii}}+\frac{d\log(d)}{T}\sqrt{\Theta_{ii}},$ (70)
$\|\Delta_{4,k}\|_{\rm S}\leq\left\|\frac{1}{T}\sum_{t=1}^{T}(f_{jt},\ j\neq i)^{\top}e_{t}^{\top}\right\|_{2}\lesssim\sqrt{\frac{d\log(d)}{T}}\sqrt{\lambda_{1}}+\frac{d\log(d)}{T}\sqrt{\lambda_{1}}.$
@@
$\|\Delta_{2,k}(g_{i\ell},g_{i\ell},\ell\neq k)\|_{\rm 2}\leq\frac{C_{1}d_{k}\log(d)}{T}+C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}+C_{1},$
$\|\Delta_{3,k}(g_{i\ell},\ell\neq k)\|_{\rm 2}\leq C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\Theta_{ii}}+\frac{C_{1}d_{k}\log(d)}{T}\sqrt{\Theta_{ii}},$ (71)
$\|\Delta_{4,k}(g_{i\ell},\ell\neq k)\|_{\rm 2}\leq C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\lambda_{1}}+\frac{C_{1}d_{k}\log(d)}{T}\sqrt{\lambda_{1}}.$
@@
$R_{k,i}$ $=\sqrt{\frac{d_{k}\log(d)}{T}}+\frac{d_{k}\log(d)}{T}+1+\sqrt{\frac{\Theta_{ii}d_{k}\log(d)}{T}}+\frac{\sqrt{\Theta_{ii}}d_{k}\log(d)}{T},\ R^{*}=\max_{i}\max_{k}R_{k,i}/\Theta_{ii}.$ (72)
@@
where the last inequality comes from condition (23) with $\rho<1$. As $R^{*}\lesssim 1$, we have $R^{*}\lesssim\psi^{\text{\footnotesize ideal}}$. It follows that, after $O(\log(\psi_{0}/\psi^{\text{\footnotesize ideal}}))$ iterations,
@@
Note that $\mathbb{E}\|e_{t,ik}\|_{2}^{2}=\mathbb{E}\xi_{t}^{\top}H^{\top}P_{g_{ik}}^{\top}P_{g_{ik}}H\xi_{t}=\mbox{tr}(HH^{\top}P_{g_{ik}}^{\top}P_{g_{ik}})\asymp d_{k}$, and $\|H^{\top}P_{g_{ik}}^{\top}P_{g_{ik}}H\|_{\rm F}^{2}=\|HH^{\top}P_{g_{ik}}^{\top}P_{g_{ik}}\|_{\rm F}^{2}\asymp d_{k}$. Choosing $x\asymp d_{k}$, we have
$\mathbb{P}\left(\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\right)\leq 4\exp\left(-C^{\prime}d_{k}^{\frac{\vartheta}{2\vartheta+4}}\right).$
Let $N:=\|e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\|_{2}$ and $\sigma_{0}^{2}:=\|\sum_{t=1}^{T}\mathbb{E}(e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}})^{2}\|_{2}$. Then, by Assumption 4.1, $N\leq C^{2}d_{k}$ and $\sigma_{0}^{2}\leq C_{0}Td_{k}$. By matrix Bernstein inequality (see, e.g., Theorem 5.4.1 of Vershynin, (2018)),
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}\left[e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right]\right\|_{2}\geq x\right)\leq 2d_{k}\exp\left(-\frac{x^{2}/2}{\sigma_{0}^{2}+Nx/3}\right).$
Choosing $x\asymp\sqrt{Td_{k}\log(d)}+d_{k}\log(d)$, with probability at least $1-d^{-c_{1}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2}\leq C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}+C_{1}\cdot\frac{d_{k}\log(d)}{T}$ (76)
Define $M:=\{1\leq t\leq T:\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}$. Since $\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}$ are independent Bernoulli random variable and $\log(T)\leq d_{k}^{\vartheta/(2\vartheta+4)}$, we have
$\mathbb{E}|M|=T\mathbb{P}\left(\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\right)\leq 4T\exp\left(-C^{\prime}d_{k}^{\frac{\vartheta}{2\vartheta+4}}\right)\leq T^{-c_{2}}.$
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}\geq x\right)$ $\leq\mathbb{P}\left(|M|\max_{t}\|e_{t,ik}\|_{2}^{2}\geq x\right)$
@@
Choosing $x\asymp d_{k}$, we have, with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}\leq C_{2}\cdot\frac{d_{k}}{T}.$ (77)
@@
$\mathbb{P}\left(\left\|\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}>0\right)=\mathbb{P}\left(\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\right)\leq 4\exp\left(-C^{\prime}d_{k}^{\frac{\vartheta}{2\vartheta+4}}\right)\leq T^{-c_{3}}.$ (78)
Combing (76), (77), (78), in an event with probability at least $1-T^{-c_{4}}-d^{-c_{1}}$,
@@
$\leq\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2}+\|P_{g_{ik}}HH^{\top}P_{g_{ik}}^{\top}\|_{2}$
$\quad+\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}+\left\|\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}$
$\leq C_{2}\sqrt{\frac{d_{k}\log(d)}{T}}+C_{2}\cdot\frac{d_{k}\log(d)}{T}+C_{2}$
@@
$N_{1}:=\left\|f_{it}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2},$
$\sigma_{1}^{2}:=\max\left\{\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}f_{it}^{2}e_{t,ik}^{\top}e_{t,ik}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2},\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}f_{it}^{2}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2}\right\}.$
@@
$N_{1}\leq C\sqrt{d_{k}}\left\|f_{it}a_{i}\right\|_{2},$
$\sigma_{1}^{2}\leq C_{3}Td_{k}\max\left\{\left\|\frac{1}{T}\sum_{t=1}^{T}f_{it}^{2}\right\|_{2},\left\|\frac{1}{Td_{k}}\sum_{t=1}^{T}f_{it}^{2}\right\|_{2}\right\}:=\sigma_{2}^{2}.$
@@
¯$\overline{\mathbb{P}}\left(\left\|\sum_{t=1}^{T}\left[f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\overline{\mathbb{E}}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right]\right\|_{2}\geq x\right)\leq 2d_{k}\exp\left(-\frac{x^{2}/2}{\sigma_{1}^{2}+N_{1}x/3}\right).$
Choosing $x\asymp\sqrt{d_{k}}\log(d)\|f_{it}\|_{2}+\sqrt{\log(d)}\sigma_{2}$, with probability at least $1-d^{-c_{4}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\overline{\mathbb{E}}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2}\leq C_{4}\frac{\sqrt{\log(d)}\sigma_{2}+\sqrt{d_{k}}\log(d)\|f_{it}a_{i}\|_{2}}{T}.$
As $\sqrt{r}\log(T)^{1/\gamma_{1}}\lesssim\sqrt{d_{k}}$, by Assumption 4.2, with probability at least $1-T^{-c_{5}}$,
$\left\|f_{it}\right\|_{2}\lesssim\sqrt{r}(\log(T))^{1/\gamma_{1}}\sqrt{\Theta_{ii}}\lesssim\sqrt{d_{k}\Theta_{ii}}.$
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\mathbb{E}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2}\leq C_{5}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\Theta_{ii}}+C_{5}\cdot\frac{d_{k}\log(d)\sqrt{\Theta_{ii}}}{T}.$

### v2

@@ Appendix A Proofs of Main Theorem
$\Delta_{3,k}(\widetilde{g}_{i\ell},\ell\neq k)=\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}{\cal E}_{t}\times_{\ell=1,\ell\neq k}^{K}\widetilde{g}_{i\ell}^{\top}\in\mathbb{R}^{d_{k}},$
$\Delta_{4,k}(\widetilde{g}_{i\ell},\ell\neq k)=\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes(w_{j}f_{jt},\ j\neq i)^{\top}\times_{\ell=1,\ell\neq k}^{K}\widetilde{g}_{i\ell}^{\top}\in\mathbb{R}^{d_{k}\times(r-1)}.$
@@
[…]
@@
$\Delta_{1,1}=a_{ik}\left(\prod_{\ell\neq k}^{K}a_{i\ell}^{\top}\widehat{g}_{i\ell}^{(m)}\right)\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{i,t}\cdot(w_{j_{2}}f_{j_{2},t},\ j_{2}\neq i)\ {\rm diag}\left(\prod_{\ell\neq k}^{K}a_{j_{2}\ell}^{\top}\widehat{g}_{i\ell}^{(m)},j_{2}\neq i\right)(a_{j_{2}k},j_{2}\neq i)^{\top}.$
@@
Let $\Upsilon_{0,i,k}=T^{-1}\sum_{t=1}^{T}w_{i}^{2}f_{it}^{2}a_{ik}a_{ik}^{\top}$ and $\Upsilon_{0,-i,k}=T^{-1}\sum_{t=1}^{T}\sum_{j_{1},j_{2}\neq i}^{r}w_{j_{1}}w_{j_{2}}f_{j_{1}t}f_{j_{2}t}a_{j_{1}k}a_{j_{2}k}^{\top}$. Then, in the event $\Omega_{1}$,
@@
$\|\Delta_{2,k}\|_{\rm S}\leq\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t}e_{t}^{\top}\right\|_{2}\lesssim\sqrt{\frac{d\log(d)}{T}}+\frac{d\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+1,$
$\|\Delta_{3,k}\|_{\rm S}\leq\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}e_{t}^{\top}\right\|_{2}\lesssim\sqrt{\frac{d\log(d)}{T}}\sqrt{\Theta_{ii}}+\frac{\sqrt{dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T}\sqrt{\Theta_{ii}},$ (71)
$\|\Delta_{4,k}\|_{\rm S}\leq\left\|\frac{1}{T}\sum_{t=1}^{T}(w_{j}f_{jt},\ j\neq i)^{\top}e_{t}^{\top}\right\|_{2}\lesssim\sqrt{\frac{d\log(d)}{T}}\sqrt{\lambda_{1}}+\frac{\sqrt{dr}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T}\sqrt{\lambda_{1}}.$
@@
$\|\Delta_{2,k}(g_{i\ell},g_{i\ell},\ell\neq k)\|_{\rm 2}\leq C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}+\frac{C_{1}d_{k}\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+C_{1},$
$\|\Delta_{3,k}(g_{i\ell},\ell\neq k)\|_{\rm 2}\leq C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\Theta_{ii}}+\frac{C_{1}\sqrt{d_{k}r}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T}\sqrt{\Theta_{ii}},$ (72)
$\|\Delta_{4,k}(g_{i\ell},\ell\neq k)\|_{\rm 2}\leq C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\lambda_{1}}+\frac{C_{1}\sqrt{d_{k}r}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T}\sqrt{\lambda_{1}}.$
@@
$R_{k,i}$ $=\sqrt{\frac{d_{k}\log(d)}{T}}+\frac{d_{k}\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+1+\sqrt{\frac{\Theta_{ii}d_{k}\log(d)}{T}}+\frac{\sqrt{\Theta_{ii}}\sqrt{d_{k}r}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}}{T},$ (73)
$R^{*}$ $=\max_{i}\max_{k}R_{k,i}/\Theta_{ii}.$ (74)
@@
where the last inequality comes from condition (23) with $\rho<1$. As $R^{*}\lesssim 1$ and $\psi_{0}\lesssim 1$, we have $R^{*}\lesssim\psi^{\text{\footnotesize ideal}}$. It follows that, after $O(\log(\psi_{0}/\psi^{\text{\footnotesize ideal}}))$ iterations,
@@
Note that $\mathbb{E}\|e_{t,ik}\|_{2}^{2}=\mathbb{E}\xi_{t}^{\top}H^{\top}P_{g_{ik}}^{\top}P_{g_{ik}}H\xi_{t}=\mbox{tr}(HH^{\top}P_{g_{ik}}^{\top}P_{g_{ik}})\asymp d_{k}$, and $\|H^{\top}P_{g_{ik}}^{\top}P_{g_{ik}}H\|_{\rm F}^{2}=\|HH^{\top}P_{g_{ik}}^{\top}P_{g_{ik}}\|_{\rm F}^{2}\asymp d_{k}$. Choosing $x\asymp d_{k}(\log T)^{(2\vartheta+4)/\vartheta}$, we have
$\mathbb{P}\left(\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\right)\leq 4\exp\left(-C^{\prime}d_{k}^{\frac{\vartheta}{2\vartheta+4}}\log T\right).$
Let $N:=\|e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\|_{2}$ and $\sigma_{0}^{2}:=\|\sum_{t=1}^{T}\mathbb{E}(e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}})^{2}\|_{2}$. Then, by Assumption 4.1, $N\leq C^{2}d_{k}(\log T)^{(2\vartheta+4)/\vartheta}$ and $\sigma_{0}^{2}\leq C_{0}Td_{k}$. By matrix Bernstein inequality (see, e.g., Theorem 5.4.1 of Vershynin, (2018)),
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}\left[e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right]\right\|_{2}\geq x\right)$
$\leq$ $2d_{k}\exp\left(-\frac{x^{2}/2}{\sigma_{0}^{2}+Nx/3}\right).$
Choosing $x\asymp\sqrt{Td_{k}\log(d)}+d_{k}\log(d)(\log T)^{(2\vartheta+4)/\vartheta}$, with probability at least $1-d^{-c_{1}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{1}\sqrt{\frac{d_{k}\log(d)}{T}}+C_{1}\cdot\frac{d_{k}\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}.$ (78)
Define $M:=\{1\leq t\leq T:\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}$. Since $\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}$ are independent Bernoulli random variable, we have
$\mathbb{E}|M|=T\mathbb{P}\left(\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{1+\frac{2}{\vartheta}}\right)\leq 4T\exp\left(-C^{\prime}d_{k}^{\frac{\vartheta}{2\vartheta+4}}\log T\right)\leq T^{-c_{2}}.$
@@
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\geq x\right)$ $\leq\mathbb{P}\left(|M|\max_{t}\|e_{t,ik}\|_{2}^{2}\geq x\right)$
@@
Choosing $x\asymp d_{k}(\log T)^{(2\vartheta+4)/\vartheta}$, we have, with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\leq C_{2}\cdot\frac{d_{k}(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}.$ (79)
@@
$\mathbb{P}\left(\left\|\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}>0\right)=\mathbb{P}\left(\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{\frac{\vartheta+2}{\vartheta}}\right)$
$\leq$ $4\exp\left(-C^{\prime}d_{k}^{\frac{\vartheta}{2\vartheta+4}}\log T\right)\leq T^{-c_{3}}.$ (80)
Combing (A), (79), (A), in an event with probability at least $1-T^{-c_{4}}-d^{-c_{1}}$,
@@
$\leq\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}+\|P_{g_{ik}}HH^{\top}P_{g_{ik}}^{\top}\|_{2}$
$\quad+\left\|\frac{1}{T}\sum_{t=1}^{T}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}+\left\|\mathbb{E}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq C_{2}\sqrt{\frac{d_{k}\log(d)}{T}}+C_{2}\cdot\frac{d_{k}\log(d)(\log T)^{\frac{2\vartheta+4}{\vartheta}}}{T}+C_{2}.$
@@
$N_{1}:=\left\|w_{i}f_{it}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2},$
$\sigma_{1}^{2}:=\max\left\{\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}w_{i}^{2}f_{it}^{2}e_{t,ik}^{\top}e_{t,ik}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2},\left\|\sum_{t=1}^{T}\overline{\mathbb{E}}w_{i}^{2}f_{it}^{2}e_{t,ik}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\right\}.$
@@
$N_{1}\leq C\sqrt{d_{k}}(\log T)^{1+\frac{2}{\vartheta}}\left\|w_{i}f_{it}a_{i}\right\|_{2},$
$\sigma_{1}^{2}\leq C_{3}Td_{k}\max\left\{\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}^{2}f_{it}^{2}\right\|_{2},\left\|\frac{1}{Td_{k}}\sum_{t=1}^{T}w_{i}^{2}f_{it}^{2}\right\|_{2}\right\}:=\sigma_{2}^{2}.$
@@
$\overline{\mathbb{P}}\left(\left\|\sum_{t=1}^{T}\left[w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\overline{\mathbb{E}}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right]\right\|_{2}\geq x\right)$
$\leq$ $2d_{k}\exp\left(-\frac{x^{2}/2}{\sigma_{1}^{2}+N_{1}x/3}\right).$
Choosing $x\asymp\sqrt{d_{k}}\log(d)(\log T)^{(\vartheta+2)/\vartheta}\|w_{i}f_{it}\|_{2}+\sqrt{\log(d)}\sigma_{2}$, with probability at least $1-d^{-c_{4}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\overline{\mathbb{E}}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{4}\frac{\sqrt{\log(d)}\sigma_{2}+\sqrt{d_{k}}\log(d)(\log T)^{(\vartheta+2)/\vartheta}\|w_{i}f_{it}a_{i}\|_{2}}{T}.$
By Assumption 4.2, with probability at least $1-T^{-c_{5}}$,
$\left\|w_{i}f_{it}\right\|_{2}\lesssim\sqrt{r}(\log(T))^{1/\gamma_{1}}\sqrt{\Theta_{ii}}.$
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{5}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\Theta_{ii}}+C_{5}\cdot\frac{\sqrt{d_{k}r}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\Theta_{ii}}}{T}.$